# The Open Journal of Astrophysics

### v1

@@ 3.1.1 Redshift evolution
As an understanding of IA is not only crucial at $z=0$, but also at higher redshifts (see Section 1), it is important to consider the redshift evolution of the variables we intend to use to select our samples. Figures 2, 3 and 4 show the variable distributions for each simulation between $0\leq z\leq 1$.
Figure 2 shows the redshift evolution for each of the variables in Figure 1, $\mathrm{BTR}$ (top left), $|v/\sigma|$ (top right), $\kappa_{\mathrm{rot}}$ (bottom left) and rir-i (bottom right), in TNG300. In agreement with observations (e.g. Maraston et al., 2009), the colours, represented by rir-i, shift to lower (bluer) values with increasing redshift and the size of the peak on the red end of the colour spectrum diminishes. This is due to the fact that the galaxy population is younger and more star-forming at higher redshift and there are less large red ellipticals, as they are formed later in the history of the universe.
@@
The distributions of the kinematic variables, $\mathrm{BTR}$, $|v/\sigma|$ and $\kappa_{\mathrm{rot}}$, show a subtle redshift evolution in TNG300. Recall from Section 2.4, that $\mathrm{BTR}$ has an inconsistent mass cut between redshifts and with the other variables. Using the transparent blue line as the distribution at $z=0$ creates mass cut consistency, i.e. the cuts at all redshifts are $M_{\star}>10^{9.8}\mathrm{M_{\odot}}/h$, within the evolution of $\mathrm{BTR}$.
@@
The kinematic variables, $|v/\sigma|$ and $\kappa_{\mathrm{rot}}$, show a clear redshift evolution in EAGLE. Contrary to TNG300, the $|v/\sigma|$ distribution skews towards lower values as we go back in time (higher redshift), indicating that there are more galaxies with a higher velocity dispersion, when compared to their rotational velocity. Therefore, Figure 3 shows an evolution from elliptical galaxies at high redshift, to disk galaxies at low redshift. The $\kappa_{\mathrm{rot}}$ distribution evolves in a similar way: the peak skews towards lower values of $\kappa_{\mathrm{rot}}$ at higher redshift, which are generally associated with elliptical galaxies.
@@
Figure 5: Galaxy stellar mass distributions at $z=0$ (top) and $z=1$ (bottom) of the full galaxy sample (black) and elliptical (continuous line style) and disk (dashed) samples, defined based on a selection in $\mathrm{BTR}$ (medium blue), $|v/\sigma|$ (light blue), $\kappa_{\mathrm{rot}}$ (dark blue) or rir-i (orange), in TNG300 (left), Horizon-AGN (middle) and EAGLE (right). The choice of variable used in the selection of the elliptical sample has a large impact on the resulting galaxy stellar mass distribution of the sample. The stellar mass distributions of the selections evolve with redshift.
@@
3.1.2 Defining samples
The distributions of the variables that we use to perform the splits into disk and elliptical galaxy samples vary between simulations (Fig. 1 - 4). Therefore, we opt for the abundance matching method to define our samples, i.e. using a set elliptical fraction of $10\%$ of the total sample of galaxies, as simulations tend to produce less large ellipticals than found in observations (Haslbauer et al., 2022). In line with the consistent stellar mass cut for the samples across redshift, we also fix the elliptical fraction, although we know observations have found both parameters to evolve with redshift (Mortlock et al., 2013). Probing the redshift evolution of a consistent sample of galaxies, e.g. using merger trees, is outside the scope of this work.
Figure 5 shows the normalised galaxy stellar mass distributions of the full sample of galaxies (black lines) and the samples of ellipticals (continuous lines) and disks (dashed lines), when defined using $\mathrm{BTR}$ (medium blue), $|v/\sigma|$ (light blue), $\kappa_{\mathrm{rot}}$ (dark blue) or rir-i (orange), at $z=0$ (top) and $z=1$ (bottom) in TNG300 (left), Horizon-AGN (middle) and EAGLE (right). While there are some discrepancies between the simulations, the stellar mass distributions of the full samples agree well. The mass distributions of the disk samples of every morphological definition match the full sample distributions closely in all shown cases, as they contain $90\%$ of the galaxies of the full sample.
@@ 3.1.2 Defining samples
Figure 6: The quadrupoles of ellipticals (continuous lines) and disks (dashed) around all resolved galaxies at $z=0$ in TNG300 (top left), Horizon-AGN (top right) and EAGLE (bottom left). The shape samples are defined by the values of the $|v/\sigma|$ (light blue), $\kappa_{\mathrm{rot}}$ (dark blue), rir-i (orange) or $M_{\star}$ (green) of the galaxies. The shapes are calculated using the simple inertia tensor. The simulations agree on the main trends, and diverge on details.
@@ 3.2 Disks and ellipticals around galaxies
In this section, the IA quadrupoles of disk and elliptical shape samples, as defined previously, around the positions of all resolved galaxies for TNG300, EAGLE and Horizon-AGN are presented. Furthermore, in TNG300, the differences between using simple and reduced shapes are explored at two redshifts, $z=0$ and $z=1$.
Figure 6 shows the quadrupoles for disks (dashed lines) and ellipticals (continuous lines) around all resolved galaxies for TNG300 (top left panel), Horizon-AGN (top right panel) and EAGLE (bottom left panel) at $z=0$. The galaxy shapes are defined by the simple inertia tensor as described in Section 2.2. The shape sample is split into disks and ellipticals as described previously, according to rir-i colour (orange), $|v/\sigma|$ (light blue), the $\kappa_{\mathrm{rot}}$ (dark blue) or stellar mass ($M_{\star}$, dark green). The blue colours of the lines in the figures are chosen to reflect the grouping of kinematic variables, where the expectation is that the selections are related based on their definitions and stellar mass distributions (Section 2.4).
Figure 7: The quadrupoles of ellipticals (continuous lines) and disks (dashed) around all resolved galaxies in TNG300 at $z=0$ (left) and $z=1$ (right) for shapes calculated using the simple (top) or reduced (bottom) inertia tensor. The shape samples are defined by the values of the $|v/\sigma|$ (light blue), $\kappa_{\mathrm{rot}}$ (dark blue), rir-i (orange) or $M_{\star}$ (green) of the galaxies. The main trends persist across redshift and shape definition.
@@
The variables being used to define the disk and elliptical galaxy samples are all strongly correlated with the stellar mass of the galaxies. As we know from previous works (e.g. Chisari et al., 2015; Tenneti et al., 2016; Samuroff et al., 2021), stellar mass also has a strong correlation to the signal strength of the galaxy alignment correlation functions. The morphological definition used to select ellipticals leads to different signal amplitudes in TNG300, that can largely be traced back to their underlying stellar mass distributions (Fig. 5), as also supported by the amplitudes of the stellar mass split (green). At $r\gtrsim 2\mathrm{Mpc}/h$, the IA signal amplitude follows the trend where more high-mass galaxies in the sample corresponds to a higher amplitude. At $r\lesssim 2\mathrm{Mpc}/h$, this is also largely the case, as the amplitude of the rir-i elliptical signal is higher than those of $|v/\sigma|$ and $\kappa_{\mathrm{rot}}$. However, at these smaller scales, we are inside the haloes and sub-grid physics models play a larger role, which is why the rir-i elliptical quadrupoles overlap with the $M_{\star}$ ones despite their diverging mass distributions; and the $|v/\sigma|$ and $\kappa_{\mathrm{rot}}$ quadrupoles are distinct, despite their overlapping mass distributions. Therefore, at these non-linear scales, we can conclude that not everything is determined by mass. This is further explored in Section 3.5.
@@
Comparing the ellipticals around galaxies correlations between top and the bottom left panel in Figure 7, i.e. shapes measured by the simple versus the reduced inertia tensor (Eq. 2), we see that the most notable difference is the distinction between the high stellar mass sample (dark green) and the rir-i elliptical sample (orange), at small scales ($r\lesssim 1\mathrm{Mpc}/h$). These two signals were not distinct for simple shapes, but for the reduced shapes the amplitude of the high stellar mass sample is clearly higher, indicating a possible difference between the influence of mass and colour on the IA in inner and outer regions of galaxies.
@@
Figure 8: The quadrupoles of disks around ellipticals at $z=0$ in TNG300 (top left), Horizon-AGN (top right) and EAGLE (bottom left). The disk and elliptical samples are defined by the values of the $|v/\sigma|$ (light blue), $\kappa_{\mathrm{rot}}$ (dark blue), rir-i (orange) or $M_{\star}$ (green) of the galaxies. The shapes are calculated using the simple inertia tensor. For this type of alignment measurement, the simulations show varying results.
@@ 3.3 Disks around ellipticals
Figure 8 shows the quadrupoles ($\times r^{2}$) for the shapes of disks around the positions of ellipticals, in TNG300 (top left panel), Horizon-AGN (top right) and EAGLE (bottom left) at $z=0$. The shapes are measured using the simple inertia tensor. Similar to the signals of disks and ellipticals around galaxies (Figure 6), the simulations agree with each other on the main trends of the disks around ellipticals correlations in Figure 8: the main $r$-dependence of the quadrupoles is quadratic and the signals are positive. While $r^{2}\tilde{\xi}_{g+}$ is flat in EAGLE and for most TNG300 morphological definitions, there is a residual scale dependence for the disks selected by rir-i around ellipticals correlation in TNG300 and for all morphological definitions in Horizon-AGN. The rir-i correlation in TNG300 shows a peak is at $r\sim 1$ and the Horizon-AGN $r^{2}\tilde{\xi}_{g+}$ measurements decrease with $r$. For EAGLE (bottom left panel) there is no distinction between the choices in morphological definition as all the signals agree within error bars. In Horizon-AGN (top right panel), there is no clear distinction between morphological definitions either, contrary to the ellipticals around galaxies signal shown in Fig. 6.
@@
Figure 9 displays the quadrupoles of disks around ellipticals as presented in Figure 8 at $z=1$. All colours are the same as previous figures, and again we use the simple inertia tensor to measure the galaxy shapes. Note that the mass cuts used in the selection of galaxies remain constant with redshift and therefore, Figure 9 represents the alignments of a different population of galaxies compared to Figure 8. While the measurements of $r^{2}\tilde{\xi}_{g+,2}$ in EAGLE are flat again at $z=1$, TNG300 has a scale dependence that decreases with $r$ in the 1-halo regime and increases with $r$ in the 2-halo regime. Like TNG300, Horizon-AGN has a diminished amplitude at $z=1$, which results in flat $r^{2}\tilde{\xi}_{g+,2}$ measurements, consistent with null at most scales.
Comparing the top let panels (TNG300) of Figures 9 and 8, reveals a reduced offset of the rir-i (orange) disks around ellipticals correlation to the $M_{\star}$ correlation (green) at $r\approx 0.7\mathrm{Mpc}/h$, from $\sim 2.5$ at $z=0$ to $\sim 1.7$ at $z=1$.
In Horizon-AGN, the $|v/\sigma|$ disks around ellipticals correlation (light blue) starts to reveal a negative at large scales, $r\gtrsim 5\mathrm{Mpc}/h$. To fully uncover for which types of measurements the correlation is below zero in Horizon-AGN, we explore both disk around ellipticals and disks around galaxies correlations for simple and reduced shapes at $z=0$ and $z=1$ in the following section.
@@ 3.4 Disks in Horizon-AGN
Figure 10 depicts the $r^{2}\tilde{\xi}_{g+,2}$ measurements of disks around ellipticals (continuous lines, as in Section 3.3) and disks around all galaxies (dashed lines, as in Section 3.2) in Horizon-AGN. These measurements are shown for simple (top row, defined by Eq. (1)) and reduced shapes (bottom row, defined by Eq. (2)) at $z=0$ (left column) and $z=1$ (right column); for the morphological samples as described in previous sections. As noted previously, the alignment signal amplitude in Horizon-AGN is not very strong, particularly for disks. Therefore in Figure 10, only the measurement for simple shapes at $z=0$ (top left panel) gives a clear non-zero signal for both types of measurements. The amplitude diminishes for reduced shapes compared to simple shapes and for $z=1$ when compared to $z=0$.
@@ 3.5 Mass relations
When selecting a (shape) sample based on a certain variable (e.g. $|v/\sigma|$), it is very difficult to distinguish between the effects of the variable itself, and the effects of its underlying stellar mass distribution on IA. Based on the assumption that stellar mass correlates heavily with IA amplitude, as shown in many other studies (e.g. Chisari et al., 2015; Tenneti et al., 2016; Samuroff et al., 2021), the previous sections all support the conclusion that the main trends in amplitude are due to the underlying stellar mass distributions of the sample selections, and secondary effects on small scales are due to (sub-grid) physics. This section attempts to disentangle these effects by re-weighting the elliptical shape samples in TNG300 defined by $|v/\sigma|$ and $\kappa_{\mathrm{rot}}$ so that their re-weighted stellar mass distributions fit the stellar mass distribution of the rir-i elliptical sample. The weights, $w_{j}$ in Eq. (5), are defined as follows:
@@
Recall from Figure 5, that the stellar mass distributions of the $|v/\sigma|$ and $\kappa_{\mathrm{rot}}$ ellipticals are very similar and include many low-mass galaxies, whereas the stellar mass distribution of the rir-i elliptical sample includes mainly high-mass galaxies. Due to how the samples are defined (see Section 3.1.2), all three elliptical samples include the same number of galaxies.
Figure 11 shows the quadrupoles of ellipticals around all galaxies in TNG300 for simple (top panels) and reduced (bottom panels) shapes at $z=0$ (left) and $z=1$ (right), as in Figure 7. The IA correlations of the elliptical shape samples defined by $|v/\sigma|$ (light blue), $\kappa_{\mathrm{rot}}$ (dark blue) or rir-i (orange) are shown in dashed lines. The continuous lines show the IA correlation functions of the $|v/\sigma|$ and $\kappa_{\mathrm{rot}}$ elliptical samples, when re-weighted according to equation 22.
The effect of re-weighting is different in the 1-halo and 2-halo regimes. At large scales, $r\gtrsim 3\mathrm{Mpc}/h$, the re-weighting has little effect. On small scales, $r\lesssim 3\mathrm{Mpc}/h$, there is an offset between the original correlations (dashed) and the re-weighted measurements (continuos). As the offset between the rir-i ellipticals around galaxies correlation and the $|v/\sigma|$ and $\kappa_{\mathrm{rot}}$ correlations increases when the latter two are re-weighted, the underlying stellar mass distribution can be ruled out as the main driver of IA at these non-linear scales. Furthermore, the decreased amplitude of the re-weighted correlations indicates that much of the power in the $|v/\sigma|$ and $\kappa_{\mathrm{rot}}$ elliptical samples comes from the low-mass galaxies ($M\star\lesssim 10^{10}\mathrm{M_{\odot}}/h$). To explore the contribution of different types of galaxies to the IA correlations further, a larger simulation would be needed, allowing for sub-selections in the samples.

### v2

@@ 3.1.1 Redshift evolution
As an understanding of IA is not only crucial at $z=0$, but also at higher redshifts (see Section 1), it is important to consider the redshift evolution of the variables we intend to use to select our samples. Figures 2, 3 and 4 show the variable distributions for each simulation between $0\leq z\leq 1$. For the measurement of the intrinsic alignment correlations we show the results at $z=0$ and $z=1$, to probe the redshift evolution and because the redshift distribution of source galaxies in Stage IV surveys is estimated to peak at $z\sim 1$.
Figure 2 shows the redshift evolution for each of the variables in Figure 1, $\mathrm{BTR}$ (top left), $|v/\sigma|$ (top right), $\kappa_{\mathrm{rot}}$ (bottom left) and rir-i (bottom right), in TNG300. For $z=0$ and $z=1$ the values used to split between disks and ellipticals are depicted in vertical dashed lines. In agreement with observations (e.g. Maraston et al., 2009), the colours, represented by rir-i, shift to lower (bluer) values with increasing redshift and the size of the peak on the red end of the colour spectrum diminishes. This is due to the fact that the galaxy population is younger and more star-forming at higher redshift and there are less large red ellipticals, as they are formed later in the history of the universe.
@@
The distributions of the kinematic variables, $\mathrm{BTR}$, $|v/\sigma|$ and $\kappa_{\mathrm{rot}}$, show a subtle redshift evolution in TNG300. Recall from Section 2.4, that the $\mathrm{BTR}$ catalogue is only available for $M_{\star}>10^{9.8}\mathrm{M_{\odot}}/h$ at $z>0$ and $M_{\star}>10^{9.3}\mathrm{M_{\odot}}/h$ at $z=0$, so the sample in this panel is different from those in the other panels in the figure. The transparent blue line shows the distribution of $\mathrm{BTR}$ for $M_{\star}>10^{9.8}\mathrm{M_{\odot}}/h$ at $z=0$ for internal comparison and consistency.
@@
The kinematic variables, $|v/\sigma|$ and $\kappa_{\mathrm{rot}}$, show a clear redshift evolution in EAGLE. Contrary to TNG300, the $|v/\sigma|$ distribution skews towards lower values as we go back in time (higher redshift), indicating that there are more galaxies with a higher velocity dispersion, when compared to their rotational velocity. Therefore, Figure 3 shows an evolution from elliptical galaxies at high redshift, to disk galaxies at low redshift. The $\kappa_{\mathrm{rot}}$ distribution evolves in a similar way: the peak skews towards lower values of $\kappa_{\mathrm{rot}}$ at higher redshift, which are generally associated with elliptical galaxies. Why these two simulations show opposing behaviour is unclear and left to future research.
@@
Figure 5: Galaxy stellar mass distributions at $z=0$ (top) and $z=1$ (bottom) of the full galaxy sample (black) and elliptical (continuous line style) and disk (dashed) samples, defined based on a selection in $\mathrm{BTR}$ (medium blue), $|v/\sigma|$ (light blue), $\kappa_{\mathrm{rot}}$ (dark blue) or rir-i (orange), in TNG300 (left), Horizon-AGN (middle) and EAGLE (right). The distributions are normalised by simulation volume and sample size, where the latter can be found in table 2. The choice of variable used in the selection of the elliptical sample has a large impact on the resulting galaxy stellar mass distribution of the sample. The stellar mass distributions of the selections evolve with redshift.
@@
3.1.2 Mass distributions
Figure 5 shows the normalised galaxy stellar mass distributions of the full sample of galaxies with well-resolved shapes (black lines) and the samples of ellipticals (continuous lines) and disks (dashed lines), when defined using $\mathrm{BTR}$ (medium blue), $|v/\sigma|$ (light blue), $\kappa_{\mathrm{rot}}$ (dark blue) or rir-i (orange), at $z=0$ (top) and $z=1$ (bottom) in TNG300 (left), Horizon-AGN (middle) and EAGLE (right). While there are some discrepancies between the simulations, the stellar mass distributions of the full samples agree well. The mass distributions of the disk samples of every morphological definition match the full sample distributions closely in all shown cases, as they contain $90\%$ of the galaxies of the full sample.
@@ 3.1.2 Mass distributions
Figure 6: The quadrupoles of ellipticals (continuous lines) and disks (dashed) around all resolved galaxies at $z=0$ in TNG300 (top left), Horizon-AGN (top right) and EAGLE (bottom left). The shape samples are defined by the values of the $|v/\sigma|$ (light blue), $\kappa_{\mathrm{rot}}$ (dark blue), rir-i (orange) or $M_{\star}$ (green) of the galaxies. The shapes are calculated using the simple inertia tensor. The simulations agree on the scale dependence and sign of the signal. Furthermore, the amplitudes of the elliptical correlations are all higher than those of the disks. The relative amplitudes of the morphological definitions differ between simulations.
@@ 3.2 Disks and ellipticals around galaxies
In this section, we present the IA quadrupoles of disk and elliptical shape samples, as defined previously, around the positions of all resolved galaxies for TNG300, EAGLE and Horizon-AGN. Furthermore, we explore the differences between using simple and reduced shapes at two redshifts, $z=0$ and $z=1$, in TNG300.
Figure 6 shows the quadrupoles for disks (dashed lines) and ellipticals (continuous lines) around all resolved galaxies for TNG300 (top left panel), Horizon-AGN (top right panel) and EAGLE (bottom left panel) at $z=0$. The galaxy shapes are defined by the simple inertia tensor as described in Section 2.2. The shape sample is split into disks and ellipticals as described previously, according to rir-i colour (orange), $|v/\sigma|$ (light blue), the $\kappa_{\mathrm{rot}}$ (dark blue) or stellar mass ($M_{\star}$, dark green), where stellar mass is not a true morphological split, but has been added as a reference.
Figure 7: The quadrupoles of ellipticals (continuous lines) and disks (dashed) around all resolved galaxies in TNG300 at $z=0$ (left) and $z=1$ (right) for shapes calculated using the simple (top) or reduced (bottom) inertia tensor. The shape samples are defined by the values of the $|v/\sigma|$ (light blue), $\kappa_{\mathrm{rot}}$ (dark blue), rir-i (orange) or $M_{\star}$ (green) of the galaxies. The scale dependence and ordering of the amplitudes of the different samples persist across redshift and shape definition indicating an underlying robustness of the signals.
@@
The variables used to define the disk and elliptical galaxy samples are all strongly correlated with the stellar mass of the galaxies. As we know from previous works (e.g. Chisari et al., 2015; Tenneti et al., 2016; Samuroff et al., 2021), stellar mass also has a strong correlation to the signal strength of the galaxy alignment correlation functions. The morphological definition used to select ellipticals leads to different signal amplitudes in TNG300, that can largely be traced back to their underlying stellar mass distributions (Fig. 5), as also supported by the amplitudes of the stellar mass split (green). At $r\gtrsim 2\mathrm{Mpc}/h$, the IA signal amplitude follows the trend where more high-mass galaxies in the sample corresponds to a higher amplitude. At $r\lesssim 2\mathrm{Mpc}/h$, this is also largely the case, as the amplitude of the rir-i elliptical signal is higher than those of $|v/\sigma|$ and $\kappa_{\mathrm{rot}}$. However, at these smaller scales, we are inside the haloes and sub-grid physics models play a larger role, which is why the rir-i elliptical quadrupoles overlap with the $M_{\star}$ ones despite their diverging mass distributions; and the $|v/\sigma|$ and $\kappa_{\mathrm{rot}}$ quadrupoles are distinct, despite their overlapping mass distributions. Therefore, at these non-linear scales, we can conclude that not everything is determined by mass. This is further explored in Section 3.5.
@@
Comparing the ellipticals around galaxies correlations between top and the bottom left panel in Figure 7, i.e. shapes measured by the simple versus the reduced inertia tensor Eq. 2, we see that the most notable difference is the distinction between the high stellar mass sample (dark green) and the rir-i elliptical sample (orange), at small scales ($r\lesssim 1\mathrm{Mpc}/h$). These two signals were not distinct for simple shapes, but for the reduced shapes the amplitude of the high stellar mass sample is clearly higher, indicating a possible difference between the influence of mass and colour on the IA in inner and outer regions of galaxies.
@@
Figure 8: The quadrupoles of disks around ellipticals at $z=0$ in TNG300 (top left), Horizon-AGN (top right) and EAGLE (bottom left). The disk and elliptical samples are defined by the values of the $|v/\sigma|$ (light blue), $\kappa_{\mathrm{rot}}$ (dark blue), rir-i (orange) or $M_{\star}$ (green) of the galaxies. The shapes are calculated using the simple inertia tensor. For this type of alignment, the simulations show differences in secondary scale-dependence and impact of morphological definition. In particular, the amplitude difference in TNG300 between the rir-i and other signals at intermediate scales indicates a fundamental difference between the two selections.
@@ 3.3 Disks around ellipticals
Figure 8 shows the quadrupoles ($\times r^{2}$) for the shapes of disks around the positions of ellipticals, in TNG300 (top left panel), Horizon-AGN (top right) and EAGLE (bottom left) at $z=0$. The shapes are measured using the simple inertia tensor. Similar to the signals of disks and ellipticals around galaxies (Figure 6), the simulations agree with each other on the main trends of the disks around ellipticals correlations in Figure 8: the main $r$-dependence of the quadrupoles is quadratic and the signals are positive. While $r^{2}\tilde{\xi}_{g+}$ does not have a strong scale dependence in EAGLE and for most TNG300 morphological definitions, there is a residual scale dependence for the disks selected by rir-i around ellipticals correlation in TNG300 and for all morphological definitions in Horizon-AGN. The rir-i correlation in TNG300 shows a peak is at $r\sim 1$ and the Horizon-AGN $r^{2}\tilde{\xi}_{g+}$ measurements decrease with $r$. For EAGLE (bottom left panel) there is no distinction between the choices in morphological definition as all the signals agree within error bars. In Horizon-AGN (top right panel), there is no clear distinction between morphological definitions either, contrary to the ellipticals around galaxies signal shown in Fig. 6.
@@
Figure 9 displays the quadrupoles of disks around ellipticals as presented in Figure 8 at $z=1$. All colours are the same as previous figures, and again we use the simple inertia tensor to measure the galaxy shapes. Note that the mass cuts used in the selection of galaxies remain constant with redshift and therefore, Figure 9 represents the alignments of a different population of galaxies compared to Figure 8. While the measurements of $r^{2}\tilde{\xi}_{g+,2}$ in EAGLE do not show a strong scale dependence at $z=1$, similar to $z=0$, TNG300 has a scale dependence that decreases with $r$ in the 1-halo regime and increases with $r$ in the 2-halo regime. Like TNG300, Horizon-AGN has a diminished amplitude at $z=1$, which results in flat $r^{2}\tilde{\xi}_{g+,2}$ measurements, consistent with null at most scales.
Comparing the top left panels (TNG300) of Figures 9 and 8, reveals a reduced offset of the rir-i (orange) disks around ellipticals correlation to the $M_{\star}$ correlation (green) at $r\approx 0.7\mathrm{Mpc}/h$, from $\sim 2.5$ at $z=0$ to $\sim 1.7$ at $z=1$.
In Horizon-AGN, the $|v/\sigma|$ disks around ellipticals correlation (light blue) starts to reveal a negative at large scales, $r\gtrsim 5\mathrm{Mpc}/h$. To fully uncover for which types of measurements the correlation is below zero in Horizon-AGN, we explore both disk around ellipticals and disks around all galaxies correlations for simple and reduced shapes at $z=0$ and $z=1$ in the following section.
@@ 3.4 Disks in Horizon-AGN
Figure 10 depicts the $r^{2}\tilde{\xi}_{g+,2}$ measurements of disks around ellipticals (continuous lines, as in Section 3.3) and disks around all galaxies (dashed lines, as in Section 3.2) in Horizon-AGN. These measurements are shown for simple (top row, defined by Eq. 1) and reduced shapes (bottom row, defined by Eq. 2) at $z=0$ (left column) and $z=1$ (right column); for the morphological samples as described in previous sections. As noted previously, the alignment signal amplitude in Horizon-AGN is not very strong, particularly for disks. Therefore in Figure 10, only the measurement for simple shapes at $z=0$ (top left panel) gives a clear non-zero signal for both types of measurements. The amplitude diminishes for reduced shapes compared to simple shapes and for $z=1$ when compared to $z=0$.
@@ 3.5 Mass relations
When selecting a (shape) sample based on a certain variable (e.g. $|v/\sigma|$), it is very difficult to distinguish between the effects of the variable itself, and the effects of its underlying stellar mass distribution on IA. Based on the assumption that stellar mass correlates heavily with IA amplitude, as shown in many other studies (e.g. Chisari et al., 2015; Tenneti et al., 2016; Samuroff et al., 2021), the previous sections all support the conclusion that the main trends in amplitude are due to the underlying stellar mass distributions of the sample selections, and secondary effects on small scales are due to (sub-grid) physics. This section attempts to disentangle these effects by re-weighting the elliptical shape samples in TNG300 defined by $|v/\sigma|$ and $\kappa_{\mathrm{rot}}$ so that their re-weighted stellar mass distributions fit the stellar mass distribution of the rir-i elliptical sample. The weights, $w_{j}$ in Eq. 5, are defined as follows:
@@
Recall from Figure 5, that the stellar mass distributions of the $|v/\sigma|$ and $\kappa_{\mathrm{rot}}$ ellipticals are very similar and include many low-mass galaxies, whereas the stellar mass distribution of the rir-i elliptical sample includes mainly high-mass galaxies. Due to how the samples are defined (see Section 3.1), all three elliptical samples include the same number of galaxies.
Figure 11 shows the quadrupoles of ellipticals around all galaxies in TNG300 for simple (top panels) and reduced (bottom panels) shapes at $z=0$ (left) and $z=1$ (right), as in Figure 7. The IA correlations of the elliptical shape samples defined by $|v/\sigma|$ (light blue), $\kappa_{\mathrm{rot}}$ (dark blue) or rir-i (orange) are shown in dashed lines. The continuous lines show the IA correlation functions of the $|v/\sigma|$ and $\kappa_{\mathrm{rot}}$ elliptical samples, when re-weighted according to Eq. 22.
The effect of re-weighting is different in the 1-halo and 2-halo regimes. At large scales, $r\gtrsim 3\mathrm{Mpc}/h$, the re-weighting has little effect. On small scales, $r\lesssim 3\mathrm{Mpc}/h$, there is an offset between the original correlations (dashed) and the re-weighted measurements (continuous). As the offset between the rir-i ellipticals around galaxies correlation and the $|v/\sigma|$ and $\kappa_{\mathrm{rot}}$ correlations increases when the latter two are re-weighted, the underlying stellar mass distribution can be ruled out as the main driver of IA at these non-linear scales. Furthermore, the decreased amplitude of the re-weighted correlations indicates that much of the power in the $|v/\sigma|$ and $\kappa_{\mathrm{rot}}$ elliptical samples comes from the low-mass galaxies ($M\star\lesssim 10^{10}\mathrm{M_{\odot}}/h$). To explore the contribution of different types of galaxies to the IA correlations further, a larger simulation would be needed, allowing for sub-selections in the samples.